# RIFT: Closed-Loop RL Fine-Tuning for Realistic and Controllable Traffic Simulation

### v1

@@ Abstract
Achieving both realism and controllability in interactive closed-loop traffic simulation remains a key challenge in autonomous driving. Data-driven simulation methods reproduce realistic trajectories but suffer from covariate shift in closed-loop deployment, compounded by simplified dynamics models that further reduce reliability. Conversely, physics-based simulation methods enhance reliable and controllable closed-loop interactions but often lack expert demonstrations, compromising realism. To address these challenges, we introduce a dual-stage AV-centered simulation framework that conducts open-loop imitation learning pre-training in a data-driven simulator to capture trajectory-level realism and multimodality, followed by closed-loop reinforcement learning fine-tuning in a physics-based simulator to enhance controllability and mitigate covariate shift. In the fine-tuning stage, we propose RIFT, a simple yet effective closed-loop RL fine-tuning strategy that preserves the trajectory-level multimodality through a GRPO-style group-relative advantage formulation, while enhancing controllability and training stability by replacing KL regularization with the dual-clip mechanism. Extensive experiments demonstrate that RIFT significantly improves the realism and controllability of generated traffic scenarios, providing a robust platform for evaluating autonomous vehicle performance in diverse and interactive scenarios. Project Page: https://currychen77.github.io/RIFT/
@@ 1 Introduction
Reliable closed-loop traffic simulation is critical for developing advanced autonomous driving systems, supporting training and evaluation [1, 2]. An ideal simulation should possess two key properties: realistic, reflecting real-world driving behavior; controllable, enabling customizable interaction styles. Existing studies typically focus either on generating realistic traffic scenarios [3, 4] or on constructing controllable interactions [5, 6, 7]. Jointly ensuring realism and controllability in interactive closed-loop scenarios remains an open problem.
Traffic simulation approaches can be broadly categorized based on the underlying simulation platform, each presenting distinct trade-offs between realism and controllability, as illustrated in Figure 1. Methods based on data-driven simulators exploit real-world data to generate realistic trajectories by learning multimodal behavioral patterns through imitation learning (IL), conditioned on the initial scene context [8, 9]. However, their open-loop training paradigm introduces the covariate shift problem during closed-loop deployment, arising from the distribution mismatch between training and deployment states. Moreover, data-driven simulators often adopt a unified vehicle dynamics model for all agents and simplified environment dynamics [10, 11], resulting in unrealistic interactions and state transitions that further degrade closed-loop reliability. In contrast, physics-based simulators provide fine-grained control over scenario configuration through physical engines, enabling high-fidelity closed-loop interactions. Nonetheless, the absence of expert demonstrations makes it challenging to reproduce realistic behavior. To mitigate this, several approaches employ reinforcement learning (RL) to acquire controllable behaviors directly through interaction with the simulator [12, 13, 14, 15], though often at the cost of realism. Other approaches enhance realism by injecting real-world traffic data into physics-based simulators [16, 17], but typically rely on log-replay or rule-based simulation, limiting controllability and interactivity. Despite recent advances, a fundamental trade-off persists between realism and controllability across both paradigms, making it challenging to achieve both simultaneously in interactive closed-loop scenarios.
@@
Drawing inspiration from the widely adopted “pre-training and fine-tuning” paradigm in large language models (LLMs) [18, 19, 20], we combine the strengths of two traffic simulation paradigms. Specifically, we perform open-loop IL pre-training in a data-driven simulator to capture trajectory-level realism and multimodality, followed by closed-loop RL fine-tuning in a physics-based simulator to address covariate shift and enhance controllability in interactive scenarios.
Building on this foundation, we introduce a dual-stage AV-centered simulation framework, as illustrated in Figure 1, which integrates the strengths of both data-driven and physics-based simulators through the “pre-training and fine-tuning” paradigm, thereby balancing realism and controllability in traffic simulation. Within this framework, critical background vehicles (CBVs) are first identified via route-level interaction analysis to construct interactive scenarios, focusing on agents most likely to interact with the AV. We then leverage an IL pre-trained trajectory generation model to produce multimodal and realistic CBV trajectories, ensuring trajectory-level realism. In the fine-tuning stage, we propose RIFT, a simple yet effective closed-loop RL fine-tuning strategy designed to enhance controllability and mitigate covariate shift. While previous approaches [21, 22] typically fine-tune only the best-performing trajectory or action, RIFT evaluates all candidate trajectories using a GRPO-style [20] group-relative advantage, preserving trajectory-level multimodality during fine-tuning. Additionally, by replacing KL regularization with a dual-clip surrogate objective, RIFT improves controllability and training stability in highly interactive closed-loop scenarios. Our contributions can be summarized as:
  • We propose a dual-stage AV-centered simulation framework that combines open-loop IL pre-training in a data-driven simulator to capture trajectory-level realism and multimodality with closed-loop RL fine-tuning in a physics-based simulator to enhance controllability and mitigate covariate shift. This decoupled paradigm leverages the complementary strengths of both platforms, balancing realism and controllability in interactive closed-loop scenarios.
  • We propose RIFT, a simple yet effective closed-loop RL fine-tuning strategy that enhances controllability and mitigates covariate shift in interactive closed-loop scenarios. During fine-tuning, RIFT maintains trajectory-level multimodality via a GRPO-style group-relative advantage formulation while enhancing controllability and training stability by replacing KL regularization with a dual-clip surrogate objective.
  • Extensive experiments demonstrate that RIFT enables the generation of realistic and controllable scenarios, and improves the reliability of AV evaluation under diverse and interactive scenarios.
@@ 2 Related Work
Realistic Traffic Simulation. A variety of generative architectures have been explored for realistic traffic simulation, including conditional variational autoencoders [23, 24, 25] and diffusion-based models [26, 27, 28]. However, maintaining long-term stability remains challenging due to the covariate shift between open-loop training and closed-loop deployment. Recent methods such as SMART [29], GUMP [30], Trajeglish [31], and MotionLM [32] address this issue by formulating traffic simulation as a next-token prediction (NTP) task, leveraging discrete action spaces to improve closed-loop robustness. Despite these advances, most approaches remain confined to data-driven simulation platforms [10, 11, 33, 34], which typically adopt a uniform vehicle dynamics model across agents and simplified environmental dynamics. Such oversimplifications limit the fidelity and reliability of long-term closed-loop interactions, especially in complex and interactive scenarios.
Controllable Traffic Simulation. To generate traffic scenarios aligned with user intent, recent studies have introduced diverse conditioning mechanisms. CTG [35] and MotionDiffuser [36] employ diffusion models conditioned on cost-based control signals. Language-conditioned methods, including CTG++[37], LCTGen[38], and ProSim [39], enable user specification through natural language prompts. Other strategies adopt guided sampling (SceneControl [40]), retrieval-based generation (RealGen [41]), or reward-driven causality modeling (CCDiff [42]). Though effective in improving controllability, these approaches are typically limited to open-loop or simplified environments, leaving the challenge of unifying realism and controllability in closed-loop simulation unresolved.
Closed-Loop Fine-Tuning. Covariate shift—the mismatch between open-loop training and closed-loop deployment—remains a key challenge for reliable long-term traffic simulation. To address this, recent work explores closed-loop fine-tuning strategies. Hybrid IL and RL methods [21, 22, 43] enhance robustness but typically fine-tune the entire model via RL, which often compromises realism due to the difficulty of designing human-aligned reward functions. Supervised fine-tuning approaches such as CAT-K [44] show strong performance but rely on expert demonstrations, limiting scalability. TrafficRLHF [45] improves alignment through reinforcement learning with human feedback (RLHF), but demands costly human input and suffers from reward model instability. Moreover, most existing methods focus on optimizing the best action or trajectory, ignoring the inherent multimodality of traffic simulation, thus limiting behavioral diversity during fine-tuning.
@@ 3.1 Task Redefinition
Following the widely adopted paradigm for closed-loop evaluation and training in autonomous driving [46, 47], our simulation framework includes a single autonomous vehicle (AV) navigating a predefined global route, accompanied by multiple rule-based background vehicles (BVs), forming an AV-centered closed-loop simulation environment. These BVs either provide diverse interactive data for training or serve to evaluate the AV’s robustness. Building upon this setup, we identify a subset of critical background vehicles (CBVs) that exhibit a higher likelihood of interacting with the AV. For these CBVs, the rule-based control is replaced with a well-trained trajectory generation model, enabling the synthesis of realistic and controllable behaviors in interactive closed-loop scenarios.
3.2 CBV-Centered Realistic Trajectory Generation
With recent advances in imitation learning, data-driven approaches have demonstrated strong performance in generating realistic, multimodal trajectories [48, 49, 50, 51, 52]. In fully observable simulation environments, methods like Pluto [53] can generate more reliable trajectories by leveraging ground-truth states. Moreover, its modular encoder-decoder architecture also facilitates downstream fine-tuning. Based on these advantages, we adopt Pluto as our trajectory generation model for synthesizing realistic, multimodal trajectories. Implementation details are presented below.
CBV-Centered Scene Encoding. Following [53], for each CBV in the scene, we extract its current feature $F_{\mathrm{cbv}}$, the historical features of neighboring vehicles $F_{\mathrm{neighbor}}$, and vectorized map features $F_{\mathrm{map}}$. These features are encoded into $E_{\mathrm{cbv}}\in\mathbb{R}^{1\times D}$, $E_{\mathrm{neighbor}}\in\mathbb{R}^{N_{\mathrm{neighbor}}\times D}$, and $E_{\mathrm{map}}\in\mathbb{R}^{N_{\mathrm{map}}\times D}$, respectively, where $N_{\mathrm{neighbor}}$ and $N_{\mathrm{map}}$ denote the number of neighboring vehicles and map elements, and $D$ is the embedding dimension. To model the interactions among these embeddings, we concatenate them and apply a global positional embedding ($\mathrm{PE}$) to obtain the unified scene embedding $E_{s}\in\mathbb{R}^{(1+N_{\mathrm{neighbor}}+N_{\mathrm{map}})\times D}$ as:
@@ 3.2 CBV-Centered Realistic Trajectory Generation
This scene embedding $E_{s}$ is then passed through $N$ Transformer encoder blocks for feature aggregation, yielding the final CBV-centered scene embedding $E_{\mathrm{enc}}$. Each encoder block follows the standard Transformer formulation. Specifically, the $i$-th block is defined as:
@@
Multimodal Trajectory Decoding. To capture the multimodal nature of real-world driving behaviors, we adopt the longitudinal-lateral decoupling mechanism proposed in [53]. This approach leverages reference line information to construct high-level lateral queries $Q_{\mathrm{lat}}\in\mathbb{R}^{N_{\mathrm{ref}}\times D}$, and introduces learnable longitudinal queries $Q_{\mathrm{lon}}\in\mathbb{R}^{N_{\mathrm{lon}}\times D}$. These are concatenated and projected to form the multimodal navigation query $Q_{\mathrm{nav}}\in\mathbb{R}^{N_{\mathrm{ref}}\times{N_{\mathrm{lon}}}\times D}$ as:
@@
$\mathrm{SelfAttn}$, $\mathrm{CrossAttn}$ denote multi-head self-attention and cross-attention operations, respectively. Given the decoder’s final output $Q_{\mathrm{dec}}$, two MLP heads are applied to produce the CBV-centered multimodal trajectories $\mathcal{T}\in\mathbb{R}^{N_{\mathrm{ref}}\times N_{\mathrm{lon}}\times T\times 6}$ and their corresponding confidence scores $\mathcal{S}\in\mathbb{R}^{N_{\mathrm{ref}}\times N_{\mathrm{lon}}}$:
@@
where $T$ is the prediction horizon, and each trajectory point $\tau_{t}^{i}$ encodes $[p_{x},p_{y},\cos{\theta},\sin{\theta},v_{x},v_{y}]$.
@@ 4 Methodology
Leveraging the trajectory generation model described in Section 3.2, realistic and multimodal trajectories can be generated across diverse scenarios. However, the open-loop training paradigm leaves the policy vulnerable to covariate shift, a distribution mismatch between expert states observed during training and those encountered at deployment, even equipped with techniques such as contrastive learning [54, 55] or data augmentation [53]. To mitigate this issue, we propose RIFT, a simple yet effective closed-loop RL fine-tuning strategy designed to enhance controllability and mitigate covariate shift while maintaining trajectory-level realism and multimodality. The following sections detail RIFT’s implementation within the dual-stage AV-centered simulation framework.
Figure 2: Overview of the RIFT: The upper section illustrates the overall architecture of RIFT. To enhance controllability, only the trajectory scoring head is fine-tuned, with the rest of the pre-trained network kept frozen to preserve trajectory-level realism. (a) The CBV identification mechanism introduces route-level interactions between the AV and CBVs. (b) Closed-loop fine-tuning improves user-aligned controllability and mitigates covariate shift.
4.1 Critical Background Vehicle Identification
Following [1], we address the “curse of rarity” [56] by selectively intervening in a subset of critical background vehicles (CBVs) at key interaction moments, while retaining rule-based control for non-critical agents to reduce computational overhead.
To facilitate meaningful interactions between the AV and surrounding BVs, we adopt the principle of decoupling route-level interaction from behavior-level realism, as inspired by [57]. As illustrated in Figure 2, we estimate the route-level interaction probability between the AV’s predefined global route and the candidate routes of other BVs. The vehicle with the highest interaction probability is selected as the CBV, and its route is then synthesized based on the predicted conflict point to support further AV-centered closed-loop simulation. Further identification details are provided in Section B.2.

### v2

@@ Abstract
Achieving both realism and controllability in closed-loop traffic simulation remains a key challenge in autonomous driving. Dataset-based methods reproduce realistic trajectories but suffer from covariate shift in closed-loop deployment, compounded by simplified dynamics models that further reduce reliability. Conversely, physics-based simulation methods enhance reliable and controllable closed-loop interactions but often lack expert demonstrations, compromising realism. To address these challenges, we introduce a dual-stage AV-centric simulation framework that conducts open-loop imitation learning pre-training in a data-driven simulator to capture trajectory-level realism and route-level controllability, followed by closed-loop reinforcement learning fine-tuning in a physics-based simulator to enhance style-level controllability and mitigate covariate shift. In the fine-tuning stage, we propose RIFT, a novel RL fine-tuning strategy that evaluates all candidate modalities through group-relative optimization with a dual-clip surrogate objective, enhancing style-level controllability and mitigating covariate shift, while preserving the trajectory-level realism and route-level controllability inherited from IL pre-training. Extensive experiments demonstrate that RIFT improves realism and controllability in traffic simulation while simultaneously exposing the limitations of modern AV systems in closed-loop evaluation. Project Page: https://currychen77.github.io/RIFT/
@@ 1 Introduction
Reliable closed-loop traffic simulation is critical for developing advanced autonomous vehicle (AV) systems, supporting training and evaluation feng2023d2rl ; ding2023survey . An ideal traffic simulation should possess two key properties: realistic, reflecting real-world driving behavior; controllable, enabling customizable traffic simulation according to user requirements.
To balance these two essential properties, existing traffic simulation methods adopt different trade-offs depending on the underlying platform, often favoring either realism or controllability, as illustrated in Figure˜1. Methods based on data-driven simulators exploit real-world data to generate realistic trajectories by learning multimodal behavioral patterns through imitation learning (IL) ngiam2021scene ; sun2022intersim ; feng2023trafficgen ; mahjourian2024unigen . In addition to realism, recent studies on data-driven simulators have pursued controllability by conditioning scenario generation on user-specified inputs—such as text conditions zhang2024chatscene ; tan2023LCTGen , goal conditions tan2024prosim ; rowe2024ctrlsim , or cost functions zhong2023CTG ; jiang2023motiondiffuser ; zhong2023CTG++ —producing scenarios that are both realistic and aligned with user requirements. However, their open-loop training paradigm introduces the covariate shift problem during closed-loop deployment, arising from the distribution mismatch between training and deployment states. Moreover, data-driven simulators often adopt simplified environment dynamics gulino2023waymax ; caesar2021nuplan , resulting in unrealistic interactions and state transitions that further degrade closed-loop reliability. In contrast, physics-based simulators provide fine-grained control over scenario configuration through physical engines, enabling high-fidelity closed-loop interactions. Nonetheless, the absence of expert demonstrations makes it challenging to reproduce realistic behavior. To mitigate this, several approaches employ reinforcement learning (RL) to directly acquire controllable behaviors through interaction with the simulator ding2021multimodal ; king ; chen2024frea ; zhang2023cat , although often at the cost of realism. Other approaches enhance realism by injecting real-world traffic data into physics-based simulators osinski2020carlareal ; li2023scenarionet , but typically rely on log-replay or rule-based simulation, limiting controllability and interactivity. Despite recent advances, a fundamental trade-off persists between realism and controllability across both paradigms, making it challenging to achieve both simultaneously in interactive closed-loop scenarios.
@@
Drawing inspiration from the widely adopted “pre-training and fine-tuning” paradigm in large language models (LLMs) rafailov2023dpo ; yu2025dapo ; shao2024grpo , we combine the strengths of two platforms. Specifically, we perform open-loop IL pre-training in a data-driven simulator to capture realism, followed by closed-loop RL fine-tuning in a physics-based simulator to address covariate shift and enhance controllability.
Building on this insight, we propose a dual-stage AV-centric simulation framework (Figure˜1) that unifies the strengths of data-driven and physics-based simulators through a “pre-training and fine-tuning” paradigm, balancing realism and controllability in traffic simulation. In Stage 1, we pre-train a planning model via IL to generate realistic and multimodal trajectories conditioned on given route-level reference lines. This stage achieves both trajectory-level realism—capturing realistic and multimodal behavior patterns—and route-level controllability—guaranteeing compliance with prescribed reference lines. In Stage 2, we identify critical background vehicles (CBVs) through route-level interaction analysis, focusing on those most likely to interact with the AV. For these CBVs, we leverage the IL pre-trained model from Stage 1—conditioned on their route-level reference lines—to automatically generate realistic and multimodal trajectories that remain route-level controllable. On top of these generated candidates, we introduce RIFT, a novel RL fine-tuning strategy that improves controllability over driving styles and mitigates covariate shift. Unlike prior methods zhang2023rtr ; peng2024improving that fine-tune only the best trajectory or action, RIFT evaluates all candidate modalities via group-relative optimization shao2024grpo together with a dual-clip surrogate objective to enhance style-level controllability and alleviate covariate shift, while preserving the trajectory-level realism and route-level controllability established in Stage 1. Our contributions can be summarized as:
  • We propose a dual-stage AV-centric simulation framework that combines open-loop IL pre-training in a data-driven simulator and closed-loop RL fine-tuning in a physics-based simulator, leveraging their complementary strengths to balance realism and controllability.
  • We propose RIFT, a closed-loop RL fine-tuning strategy that evaluates all candidate modalities through group-relative optimization shao2024grpo with a dual-clip surrogate objective, improving style-level controllability and alleviating covariate shift, while retaining the trajectory-level realism and route-level controllability inherited from IL pre-training.
  • Extensive experiments demonstrate that RIFT enhances the realism and controllability of traffic simulation, effectively exposing the limitations of modern AV systems under closed-loop settings.
@@ 2 Related Work
Realistic Traffic Simulation. A variety of generative architectures have been explored for realistic traffic simulation tan2021scenegen ; zhang2023trafficbots ; yang_infgen , including conditional variational autoencoders trafficsim ; rempe2022strive ; xu2023bits and diffusion-based models jiang2024scenediffuser ; chitta2024sledge ; zhou2024Nexus ; Tan_SceneDiffuser++ . However, maintaining long-term stability remains challenging due to the covariate shift between open-loop training and closed-loop deployment. Recent methods such as SMART wu2024smart , GUMP hu2024gump , Trajeglish philiontrajeglish , and MotionLM seff2023motionlm address this issue by formulating traffic simulation as a next-token prediction (NTP) task, leveraging discrete action spaces to improve closed-loop robustness. Despite these advances, most approaches remain confined to data-driven simulation platforms gulino2023waymax ; caesar2021nuplan ; Dauner2024navsim ; montali2023wosac , which typically adopt simplified environment dynamics. Such oversimplifications limit the fidelity and reliability of long-term closed-loop interactions, especially in complex and interactive scenarios.
Controllable Traffic Simulation. Recent studies have introduced diverse conditioning mechanisms to generate traffic scenarios aligned with user preferences. CTG zhong2023CTG and MotionDiffuser jiang2023motiondiffuser employ diffusion models conditioned on cost-based control signals. Language-conditioned methods, including CTG++zhong2023CTG++ , LCTGentan2023LCTGen , and ProSim tan2024prosim , enable user specification through natural language prompts. Other strategies adopt guided sampling (SceneControl lu2024scenecontrol ), retrieval-based generation (RealGen ding2024realgen ), or reward-driven causality modeling (CCDiff lin2024ccdiff ). Despite improving controllability, existing approaches remain confined to open-loop settings or simplified environment dynamics, and primarily target low-level control. High-level attributes such as driving style receive limited attention, leaving the integration of realism and controllability in closed-loop simulation an open challenge.
Closed-Loop Fine-Tuning. Covariate shift—the mismatch between open-loop training and closed-loop deployment—remains a key challenge for reliable long-term traffic simulation. To address this, recent work explores closed-loop fine-tuning strategies. Hybrid IL and RL methods zhang2023rtr ; peng2024improving ; lu2023imitationisnotenough enhance robustness but typically fine-tune the entire model via RL, which often compromises realism due to the difficulty of designing human-aligned reward functions. Supervised fine-tuning approaches such as CAT-K zhang2025catk show strong performance but rely on expert demonstrations, limiting scalability. TrafficRLHF cao2024trafficrlhf improves alignment through reinforcement learning with human feedback (RLHF), but demands costly human input and suffers from reward model instability. Moreover, most existing methods focus on optimizing the best action or trajectory, ignoring the inherent multimodality of traffic simulation, thus limiting behavioral diversity during fine-tuning.
@@ 3.1 Task Redefinition
Following the widely adopted paradigm for closed-loop training and evaluation in autonomous driving jia2024bench2drive ; xu2022safebench , our simulation framework includes a single autonomous vehicle (AV) navigating a predefined global route, accompanied by multiple rule-based background vehicles (BVs), forming an AV-centric closed-loop simulation environment. These BVs either provide diverse interactive data for training or serve to evaluate the AV’s robustness. Building upon this setup, we identify a subset of critical background vehicles (CBVs) that exhibit a higher likelihood of interacting with the AV. For these CBVs, the rule-based control is replaced with a well-trained planning model, enabling the synthesis of realistic and controllable behaviors in interactive closed-loop scenarios.
3.2 CBV-Centric Realistic Trajectory Generation
With recent advances in imitation learning, data-driven approaches have demonstrated strong performance in generating realistic, multimodal trajectories zheng2025diffusionbased ; Huang_2023_GameFormer ; hu2023_uniad ; jiang2023vad ; sun2024sparsedrive . In fully observable simulation environments, Pluto cheng2024pluto produces reliable, realistic, and multimodal trajectories by leveraging ground-truth states, while enabling route-level controllability through reference line encoding. These capabilities make Pluto a suitable choice for our planning model.
CBV-Centric Scene Encoding. Following cheng2024pluto , for each CBV in the scene, we extract its current feature $F_{\mathrm{cbv}}$, the historical features of neighboring vehicles $F_{\mathrm{neighbor}}$, and vectorized map features $F_{\mathrm{map}}$. These features are encoded into $E_{\mathrm{cbv}}\in\mathbb{R}^{1\times D}$, $E_{\mathrm{neighbor}}\in\mathbb{R}^{N_{\mathrm{neighbor}}\times D}$, and $E_{\mathrm{map}}\in\mathbb{R}^{N_{\mathrm{map}}\times D}$, respectively, where $N_{\mathrm{neighbor}}$ and $N_{\mathrm{map}}$ denote the number of neighboring vehicles and map elements, and $D$ is the embedding dimension. To model the interactions among these embeddings, we concatenate them and apply a global positional embedding ($\mathrm{PE}$) to obtain the unified scene embedding $E_{s}\in\mathbb{R}^{(1+N_{\mathrm{neighbor}}+N_{\mathrm{map}})\times D}$ as:
@@ 3.2 CBV-Centric Realistic Trajectory Generation
This scene embedding $E_{s}$ is then passed through $N$ Transformer encoder blocks for feature aggregation, yielding the final CBV-centric scene embedding $E_{\mathrm{enc}}$. Each encoder block follows the standard Transformer formulation. Specifically, the $i$-th block is defined as:
@@
Multimodal Trajectory Decoding. To capture the multimodal nature of real-world driving behaviors, we adopt the longitudinal-lateral decoupling mechanism proposed in cheng2024pluto . This approach leverages reference line information to construct high-level lateral queries $Q_{\mathrm{lat}}\in\mathbb{R}^{N_{\mathrm{ref}}\times D}$, and introduces learnable longitudinal queries $Q_{\mathrm{lon}}\in\mathbb{R}^{N_{\mathrm{lon}}\times D}$. These are concatenated and projected to form the multimodal navigation query $Q_{\mathrm{nav}}\in\mathbb{R}^{N_{\mathrm{ref}}\times{N_{\mathrm{lon}}}\times D}$ as:
@@
$\mathrm{SelfAttn}$, $\mathrm{CrossAttn}$ denote multi-head self-attention and cross-attention operations, respectively. Given the decoder’s final output $Q_{\mathrm{dec}}$, two MLP heads are applied to produce the CBV-centric multimodal trajectories $\mathcal{T}\in\mathbb{R}^{N_{\mathrm{ref}}\times N_{\mathrm{lon}}\times T\times 6}$ and their corresponding confidence scores $\mathcal{S}\in\mathbb{R}^{N_{\mathrm{ref}}\times N_{\mathrm{lon}}}$:
@@
where $T$ is the prediction horizon, and each trajectory point $\tau_{t}^{i}$ encodes $p_{x},p_{y},\cos{\theta},\sin{\theta},v_{x},v_{y}]$.
@@ 4 Methodology
Leveraging the IL pre-trained planning model described in Section˜3.2, realistic and multimodal trajectories can be generated across diverse scenarios conditioned on reference lines. However, the open-loop training paradigm leaves the policy vulnerable to covariate shift, even with contrastive learning halawa2022action ; wang2023fend or data augmentation cheng2024pluto . To address this, we propose RIFT, a closed-loop RL fine-tuning strategy that enhances style-level controllability and mitigates covariate shift while preserving the trajectory-level realism and route-level controllability from pre-training. The following sections detail RIFT’s implementation within the physics-based simulator.
Figure 2: Overview of the RIFT: Building on the IL pre-trained model, RIFT first performs route-level interaction analysis to identify critical background vehicles and their associated reference lines, enabling the generation of realistic and multimodal trajectories. To isolate style-level controllability from the trajectory-level realism and route-level controllability established during pre-training, only the scoring head is fine-tuned via RIFT, with the remaining components kept frozen. Specifically, RIFT computes group-relative advantages over all candidate rollouts, promoting alignment with user-preferred styles and mitigating covariate shift through closed-loop RL fine-tuning.
4.1 Trajectory Generation for CBVs
Following feng2023d2rl , we address the “curse of rarity” liu2024curse by selectively intervening in a small set of critical background vehicles (CBVs) at key interaction moments, while keeping non-critical agents under rule-based control to reduce computational overhead. CBVs are identified via route-level interaction analysis between the AV’s predefined global route and the candidate routes of surrounding vehicles, selecting the vehicle with the highest interaction probability (details in Section˜B.2).
The corresponding route-level reference line is then used as a condition for the IL pre-trained planning model (Section˜3.2) to synthesize realistic and multimodal trajectories. For each identified CBV, the model generates $N_{\mathrm{ref}}\times N_{\mathrm{lon}}$ candidate trajectories, from which the highest-scoring one is selected for closed-loop execution. This process promotes realistic route-level interactions with the AV and enables the construction of meaningful interactive scenarios.